OpenStudy (anonymous):

In which expression should the exponents be multiplied? (53)−4 one third to the 4th times one third to the 7th 3 to the 7th over 3 to the 15th 45 ⋅ 42

OpenStudy (anonymous):

@perl can you help me please

OpenStudy (perl):

hi, i am having trouble reading that. can you take a screenshot of the question

OpenStudy (anonymous):

i do not know how i am on computer hold on let me get my tabilt

OpenStudy (perl):

thx

OpenStudy (perl):

ok we have three exponent rules. i am going to write them down

OpenStudy (anonymous):

OK

OpenStudy (perl):

$$ \Huge \rm { \\ a^m\cdot a^n = a^{m+n} \\~\\~\\(a^m)^n = a^{m\cdot n } \\~\\~\\ \frac{a^m}{a^n}= a^{m-n} }$$

OpenStudy (perl):

call these rule 1, rule 2, rule 3. rule 2 is what we want

OpenStudy (anonymous):

Ok

OpenStudy (perl):

which of the options use rule 2 ?

OpenStudy (anonymous):

A and B

OpenStudy (perl):

how does B) use rule 2 ?

OpenStudy (perl):

B) is an example of rule 1

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

So A

OpenStudy (perl):

$$ \LARGE \rm { (5^{3})^{-4} = 5^{3\times \text{-}4 } } $$